OpenStudy (anonymous):

Solve the equation 12 + 2n = 2 For n Help really needed piz

OpenStudy (bibby):

You're solving for n. Isolate n. How do you get n alone on one side of the equation?

OpenStudy (anonymous):

Go ask Ms. Kaple.

OpenStudy (anonymous):

lol shelby no i rather ask on here lol

OpenStudy (anonymous):

N=-5

OpenStudy (bibby):

I don't believe you. I want to see some work.

OpenStudy (anonymous):

I see youuuuu(: lmao message me back

OpenStudy (anonymous):

12 +2n = 2... first subtract 12 from 2, 2x = -10 then divide by 2 x = -5

OpenStudy (blues):

HaHa, spamming other peoples' questions is most definitely not funny.

OpenStudy (anonymous):

i did tho and it cant be 5 it can only be -4 4 0 1

OpenStudy (anonymous):

were in the same class.. is okay

OpenStudy (anonymous):

Will sonme help me pizzzzz

OpenStudy (bibby):

n=-5 for 12 + 2n = 2 You probably pasted the equation wrong

OpenStudy (anonymous):

ohh i ment 12 + 2n = 20 for n my bad

OpenStudy (bibby):

Isolate 2n. How do you get rid of the 12?

OpenStudy (anonymous):

well in that case n=4

OpenStudy (anonymous):

- 4 or 4

OpenStudy (bibby):

I guess the answer is to get some guy on the internet to spoonfeed you

OpenStudy (anonymous):

4

OpenStudy (anonymous):

k thankx

OpenStudy (anonymous):